# Optimas: Optimizing Compound AI Systems with Globally Aligned Local Rewards

### v2

@@ Abstract
Figure 2: Five real-world and challenging compound AI systems. The goal is to automatically optimize the configuration across a heterogeneous set of components and parameters, e.g., model parameters, prompts, model selection choice, and hyperparameters. See Appendix C for details.
@@ 1 Introduction
Modern AI systems increasingly employ compound systems that integrate multiple complex components, such as Large Language Models (LLMs), tool/function calls, and traditional machine learning models like retrievers [1, 2, 3]. These components collaborate to process heterogeneous data sources and solve complex tasks through specialized subtask allocation [4, 5, 6, 7, 8]. While compound AI systems have yielded performance advantages over monolithic models [3, 9, 10, 11, 12], they can be highly sensitive to the failure of individual components, which leads to cascading failures in the final results [13, 14, 15, 16]. For example, if an LLM misinterprets an input query, it can retrieve irrelevant or misleading information. This leads to subsequent tool calls operating on incorrect inputs, producing unreliable outputs throughout the system. Therefore, optimizing these systems as a whole is crucial for maintaining reliability and system performance (i.e., global rewards).
@@
However, optimizing these compound AI systems end-to-end is fundamentally challenging due to their non-differentiable nature. Second, it is hard to jointly optimize heterogeneous configurations (textual vs. numerical, continuous vs. discrete) from individual components, including prompts, hyperparameters, model selections, and even model weights. Third, running the entire compound AI system during optimization to achieve global reward can be costly.
@@
Previous works have largely focused on optimizing specific configurations in isolation, such as optimizing prompts through textual feedback [1, 2, 17, 18, 19] or model selection through iterative search [8, 20]. Yet these approaches can fail to capture critical bottlenecks—for example, a perfectly optimized prompt can struggle to compensate for a poorly chosen model. Furthermore, even individually well-optimized components may interact suboptimally, as outputs from one component optimized upstream might not have visibility into what a downstream component might need to improve the global reward. Moreover, such methods typically require costly system runs on many configurations, yielding low data efficiency.
@@
Present work. Here we propose Optimas (Figure 1), a unified framework for the effective and data-efficient optimization of compound AI systems. The core idea is to learn a globally aligned Local Reward Function (LRF) for each component, such that independently maximizing a component’s local reward still reliably improves the global rewards. We show that under mild conditions, our approach converges reliably, providing strong theoretical guarantees on optimization efficacy. Furthermore, since the learned LRFs can be used to optimized components locally, Optimas has higher data efficiency by avoiding extensive runs of the entire compound AI system to achieve high global reward.
@@
Specifically, each LRF estimates the contribution of its corresponding component’s output to the global reward. All LRFs are implemented using a shared LLM backbone, with separate projection heads for each component to produce component-specific rewards. We propose a lightweight adaptation mechanism using mini-batch preference data to ensure the LRFs remain aligned with the evolving system configuration (Figure 3). Leveraging its decentralized structure, Optimas applies designated optimization method to each component based on its configuration type—for example, reinforcement learning for model parameters [21, 22] or metric-guided search for prompts and hyperparameters [17, 23, 24]. Overall, Optimas iteratively updates heterogeneous configurations towards a higher global reward by using each adaptive LRF as an objective—optimizing a component to maximize its local reward, reducing the entire system runs to maintain higher data efficiency.
@@
We conduct extensive experiments to evaluate Optimas across five real-world compound systems (Figure 2), including challenging settings such as behavior-driven product recommendation and medical analysis. Optimas consistently outperforms strong baselines, achieving an average relative improvement of 11.92% with higher data efficiency. While baseline methods occasionally improve performance—for example, DSPy improves the performance on the multi-hop QA system—they may degrade performance on other tasks, such as product recommendation. In contrast, Optimas is the only method that improves performance across all five tasks, consistent with our theoretical guarantee (Section 4.4) that aligning local and global rewards enables effective optimization.
@@
Table 1: A comparison of Optimas with selected methods. Optimas optimizes compound systems with heterogeneous configurations and enables higher data efficiency with local optimization to reduce entire system runs. We prove Optimas’s convergence under mild conditions.
@@
OPRO [17]
DSPy [2]
TextGrad [1]
LLMSelector [8]
@@ 2 Related Work
Optimizing LLM single-step generation. Prior work extensively optimizes prompts for Large Language Models (LLMs) in single-step generation to improve performance [17, 18, 25, 26, 27, 28, 29], but these methods are limited in their ability to handle complex, multi-step tasks. For example, addressing complex queries often requires combining multiple components—such as LLMs, tools, and machine learning predictors—to obtain accurate predictions.
Optimizing multi-component/multi-step generation. Compound AI systems consisting of multiple components enable more complex planning and specialized processing at each task step [2, 30, 12, 3, 31]. Previous studies typically optimize different components separately, such as optimizing LLM prompts [2, 19, 1], fine-tuning model weights using supervised learning [32] or reinforcement learning [33, 34], developing model routing [8] and layer grouping[35] strategies, and selecting hyperparameters [36, 37, 38, 39]. For example, LLMSelector [8] learns a lightweight policy to select the best LLM model for a given component at each step. However, optimizing only a subset of components might overlook critical bottlenecks to global reward. In contrast, Optimas enables end-to-end optimization across all components.
Reward modeling for multi-step tasks. To provide more fine-grained supervision, recent works break down global rewards (e.g., answer accuracy) into more targeted signals (i.e., dense/process rewards). Representative approaches include leveraging or bootstrapping from human step-wise annotations [40, 41, 42]; hierarchical planning that assigns rewards to self-reflection and error correction steps [43]; using Monte Carlo Tree Search to assign credit to intermediate reasoning steps [44, 45, 46, 47, 48] or actions [49, 50]. The key insight of Optimas is to maintain one globally aligned Local Reward Function (LRF) per component within an evolving compound system. Unlike prior work focused on differentiable single policy models, Optimas adapts these LRFs to maintain local-global alignment in a non-differentiable, multi-component system, consistently improving the global reward. The local optimization also offers higher data efficiency compared to some existing approaches [1, 8, 32] that require extensive system runs. Moreover, we provide theoretical analysis to prove the convergence of our framework. In Table 1, we highlight our key contributions compared to closely related works.
@@ 3 Problem Formulation: Optimizing Compound AI Systems
Forward execution. For a given input $x$ and configuration policy $\mathbf{v}$, the system executes components in topological order over the edge set $\mathcal{E}(x)$:
$y_{k}=C_{k}\bigl{(}\{y_{i}\mid C_{i}\in pa(C_{k})\};\,\mathbf{v}_{k}\bigr{)},$ (1)
where $pa(C_{k})$ denotes all the parents of component $C_{k}$ over $\mathcal{E}(x)$. For clarity, we define the overall system as $f(x;\,\mathbf{v}):=y$, where $y$ is a collection of outputs from one or more components.
Optimization objective. Given a training dataset $\mathcal{D}$ with system inputs and a user-defined global reward function $R:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$ that evaluates the final system output, the optimization goal is to find the configuration policy $\mathbf{v}^{\star}(x)$ that maximizes the expected global reward:
@@ 4 Optimas: Globally Aligned Local Rewards for Optimization
Challenges. Directly optimizing the objective in Eq. (2) is difficult. As the configuration spaces are typically non-differentiable, gradient-based optimization cannot be used. Moreover, each policy $\mathbf{v}_{k}$ may control a different configuration type, so the joint policy $\mathbf{v}$ can span heterogeneous spaces. Therefore, previous efforts [1, 32, 8, 2] largely focus on optimizing the policy for single types of configurations, which simplifies the optimization problem; however, this also leads to suboptimal compound systems. Moreover, even with a single type of configurations, existing methods often require extensive runs on the system to search for better configurations, resulting in low data efficiency.
Key intuition. To address the above challenges, our approach (Figure 3) is to learn Local Reward Functions (LRFs) over individual components that align with the global reward, allowing independent optimization on heterogeneous components using different optimization approaches while improving the global reward. Since the optimization with LRFs is conducted locally, we reduce the number of entire system runs to achieve a high global reward.
Moreover, as the system configurations evolve over iterations, the LRFs should adapt to remain consistent with the updated configurations. Specifically, we introduce a local-global alignment property for learning LRFs (Section 4.1). To ensure alignment, Optimas employs a lightweight adaptation mechanism that updates LRFs with minimal data sampled from the system, preserving consistency with the global reward (Section 4.2). These globally aligned LRFs enable each component to be effectively optimized through local reward maximization towards global reward gains (Section 4.3).
@@ 4.1 Learning Local Reward Functions
Definition (Local Reward Function (LRF)). An LRF on component $C_{k}$ is defined as $r_{k}:(x_{k},y_{k})\rightarrow\mathbb{R}$, which evaluates the component’s output $y_{k}$ given the provided context, such as the input $x_{k}$ and an optional component task description.
Practical implementation. In practice, we implement all LRFs with a single LLM backbone $\phi$ and separate linear heads $h_{k}$, one head per component $C_{k}$ for components with non-empty configurations. The backbone encodes the concatenated text inputs $[x_{k},y_{k}]$ into an embedding, and the corresponding head projects this embedding to a scalar reward value. Therefore, each LRF is modeled by
[…]
We represent the LRFs as one multitask network to reduce computation and memory costs while maintaining the capacity to produce component-specific rewards.
@@
where $\mathbf{v}_{-k}$ denotes the configurations of all downstream components—those that receive, directly or indirectly, information originating from $C_{k}$. The expectation term, which denotes the expected global reward for each candidate output, is estimated via Monte Carlo sampling. This involves executing the downstream components with the candidate output fixed, capturing their stochasticity, and averaging the resulting global rewards from the final system outputs.
Objective of reward functions. To obtain LRFs that conform to the local-global alignment property, we train each $r_{k}$ using a pairwise log-sigmoid ranking loss:
@@
where $\mathcal{D}_{k}(\mathbf{v})$ is a preference dataset collected under the current system configuration $\mathbf{v}$.
To generate the preference dataset for each component $C_{k}$, we 1) execute the compound system up to $C_{k}$ and record the partial trajectory $\bigl{\langle}x,\,(x_{1},y_{1}),\dots,(x_{k-1},y_{\,k-1})\bigr{\rangle}$; 2) sample two candidate outputs for $C_{k}$ (e.g., via higher-temperature decoding or alternate hyperparameters); and 3) estimate their expected task metrics according to the expectation terms on the right-hand side of Eq. (4). The output with the higher expected value is labeled as $y_{k}^{+}$, and the other as $y_{k}^{-}$ in $\mathcal{D}_{k}(\mathbf{v})$.
4.2 Adaptive Local Reward Functions via Offline-Online Training
Problem: Misaligned LRFs in the evolving system. As the system configuration evolves during optimization, LRFs trained under a previous configuration $\mathbf{v}^{t}$ may become inaccurate under the updated configuration $\mathbf{v}^{t+1}$. For example, after updating a component $C_{k}$, the same upstream outputs may lead to different final outcomes, making earlier reward estimates inaccurate. Meanwhile, downstream components now receive inputs generated by the updated $C_{k}$, which may fall outside the distribution seen by their LRFs. These shifts accumulate over time, degrading the local–global alignment property (Eq. 4) that LRFs are designed to satisfy.
However, retraining all LRFs from scratch after every configuration update lowers the data efficiency due to the system runs involved to minimize Eq.4. To address this, we develop a lightweight adaptation strategy that incrementally refines the LRFs as the system evolves, maintaining alignment without full retraining.
Offline reward modeling. Given the initial system configuration $\mathbf{v}^{0}$ and input dataset $\mathcal{D}$, we first construct an offline preference dataset for each component and train its corresponding LRF to convergence. This offline phase establishes well-aligned LRFs that accurately reflect each component’s contribution to the global task metric, based on the input and output distribution of the components, providing a reliable foundation for subsequent optimization.
Online reward function adaptation. We sample a small input batch from $\mathcal{D}$ and construct a mini-batch of preference data $\mathcal{B}_{k}$ for $C_{k}$ using the procedure described in Section 4.1. Therefore, we optimize the LRF on $C_{k}$ on the objective:
$\mathcal{L}^{\text{adapt}}_{k}(\mathcal{B}_{k})=\mathcal{L}_{k}(\mathcal{B}_{k})$ (6)
This adaptation helps maintain the local–global alignment property (Eq. 4). Optionally, we can maintain a buffer of previous generated preference data for $C_{k}$ into $\mathcal{B}_{k}$ to enable stable optimization.

### v3

@@ Abstract
Figure 2: Five real-world and challenging compound AI systems. The goal is to automatically optimize the configuration across a heterogeneous set of components and parameters, e.g., model parameters, prompts, model selection choice, and hyperparameters. See Appendix D for details.
@@ 1 Introduction
Modern AI systems increasingly employ compound systems that integrate multiple complex components, such as Large Language Models (LLMs), tool/function calls, and traditional machine learning models like retrievers (Yuksekgonul et al., 2025; Khattab et al., 2023; Du et al., 2023). These components collaborate to process heterogeneous data sources and solve complex tasks through specialized subtask allocation (Zaharia et al., 2024; Zhou et al., 2025; Leike et al., 2018; Kandogan et al., 2025; Chen et al., 2025b). While compound AI systems have yielded performance advantages over monolithic models (Du et al., 2023; Huang et al., 2023; Leblond & et al., 2023; Lewis et al., 2020; Liu et al., 2023), they can be highly sensitive to the failure of individual components, which leads to cascading failures in the final results (Cemri et al., 2025; tse Huang et al., 2025; Peng et al., 2025; Chen et al., 2024a). For example, if an LLM misinterprets an input query, it can retrieve irrelevant or misleading information. This leads to subsequent tool calls operating on incorrect inputs, producing unreliable outputs throughout the system. Therefore, optimizing these systems as a whole is crucial for maintaining reliability and global system performance (i.e., global rewards).
@@
However, optimizing these compound AI systems end-to-end is fundamentally challenging due to their non-differentiable nature. Second, it is hard to jointly optimize heterogeneous configurations (textual or numerical, continuous or discrete) from individual components, including prompts, hyperparameters, model selections, and even model weights. Moreover, running the entire compound AI system during optimization to achieve global reward is costly.
@@
Previous works have largely focused on optimizing specific configurations in isolation, such as optimizing prompts through textual feedback (Yuksekgonul et al., 2025; Khattab et al., 2023; Yang et al., 2024; Madaan et al., 2023; Wu et al., 2024a) or model selection through iterative search (Chen et al., 2025b; 2024c). Yet these approaches can fail to capture critical bottlenecks. For example, a perfectly optimized prompt can struggle to compensate for a poorly chosen model. Even when the components are individually well optimized, they may still collaborate suboptimally, as the upstream component might not have visibility into which inputs are effective for the downstream components. Consequently, previous methods typically require costly system runs over many configurations to identify the best configuration for the components to work well together, leading to low data efficiency due to the large number of configurations.
@@
Present work. Here we propose Optimas (Figure 1), a unified framework for the effective and data-efficient optimization of compound AI systems. The core idea is to learn a globally aligned Local Reward Function (LRF) per component, such that independently maximizing a component’s local reward still reliably improves the global rewards. We show that under mild conditions, our approach converges reliably, providing strong theoretical guarantees. Furthermore, since the learned LRFs can be used to optimized components locally, Optimas has higher data efficiency by avoiding extensive runs of the entire compound AI system to achieve high global reward.
@@
Specifically, each LRF estimates the contribution of a component’s output to the global reward. All LRFs are implemented using a shared LLM backbone, with separate projection heads for each component to produce component-specific rewards. We propose a lightweight adaptation mechanism using mini-batch preference data to ensure the LRFs remain aligned with the evolving system configuration (Figure 3). Leveraging this decentralized structure, Optimas applies specific optimization method to each component based on its configuration type. For example, reinforcement learning for model parameters (Rafailov et al., 2023; Schulman et al., 2017) or metric-guided search for prompts and hyperparameters (Yang et al., 2024; Opsahl-Ong et al., 2024; Liashchynskyi & Liashchynskyi, 2019). Overall, Optimas iteratively updates heterogeneous configurations towards a higher global reward by using each adaptive LRF as an objective. By optimizing a component to maximize its local reward, Optimas reduces the entire system runs to maintain higher data efficiency.
@@
We conduct extensive experiments to evaluate Optimas across five real-world compound systems (Figure 2), including challenging settings such as behavior-driven product recommendation and medical analysis. Optimas consistently outperforms strong baselines, achieving an average relative improvement of 11.92% with higher data efficiency. While baseline methods occasionally improve performance. For example, while DSPy improves the performance on the multi-hop QA system, it may degrade performance on other tasks, such as product recommendation. In contrast, Optimas is the only method that improves performance across all five tasks, consistent with our theoretical guarantee (Section 4.4) that aligning local and global rewards enables effective optimization.
@@
Table 1: A comparison of Optimas with selected methods. Optimas optimizes compound systems with heterogeneous configurations and enables higher data efficiency with local optimization to reduce number of system runs. We prove Optimas’s convergence under mild conditions.
@@
OPRO (Yang et al., 2024)
DSPy (Khattab et al., 2023)
TextGrad (Yuksekgonul et al., 2025)
LLMSelector (Chen et al., 2025b)
@@ 2 Related Work
Optimizing LLM single-step generation. Prior work extensively optimizes prompts for Large Language Models (LLMs) in single-step generation to improve performance (Yang et al., 2024; Madaan et al., 2023; Guo et al., 2024; Shinn et al., 2023; Yu et al., 2024; Yin et al., 2025; Chen et al., 2025a), but these methods are limited in their ability to handle complex, multi-step tasks. For example, addressing complex queries often requires combining multiple components—such as LLMs, tools, and machine learning predictors—to obtain accurate predictions.
Optimizing multi-component/multi-step generation. Compound AI systems consisting of multiple components enable more complex planning and specialized processing at each task step (Khattab et al., 2023; Yao et al., 2023; Liu et al., 2023; Du et al., 2023; Zhang et al., 2025). Previous studies typically optimize different components separately, such as optimizing LLM prompts (Khattab et al., 2023; Wu et al., 2024a; Yuksekgonul et al., 2025), fine-tuning model weights using supervised learning (Zhao et al., 2025; Chen et al., 2024d) or reinforcement learning (Lin et al., 2023; Chen et al., 2025c), developing model routing (Chen et al., 2025b) and layer grouping (Chen et al., 2023) strategies, and selecting hyperparameters (Wang et al., 2023; Falkner et al., 2018; Pham et al., 2018; Liu et al., 2019). In contrast, Optimas enables end-to-end optimization across all components.
Reward modeling for multi-step tasks. To provide more fine-grained supervision, recent works break down global rewards (e.g., answer accuracy) into more targeted signals (i.e., dense/process rewards). Representative approaches include leveraging or bootstrapping from human step-wise annotations (Lightman et al., 2024; Uesato et al., 2022; She et al., 2025); hierarchical planning that assigns rewards to error correction steps (Wang et al., 2025); using Monte Carlo Tree Search to assign credit to intermediate reasoning steps (Wang et al., 2024; Ma et al., 2025; Jiao et al., 2024; Chen et al., 2024b; Setlur et al., 2025) or actions (Chen et al., 2025d; Choudhury, 2025). Recently, Chen et al. (2024d) leverage Bayesian optimization to decompose global losses into local losses for optimizing model weight. Optimas differs by dynamically aligning local rewards with global rewards through preference-based adaptation. This design is applicable to both differentiable and non-differentiable configurations, without requiring fixed decomposition or extensive retraining. The local optimization avoids extensive system runs and offers higher data efficiency. Moreover, we provide theoretical analysis to prove the convergence of our framework. We highlight our key contributions in Table 1.
@@ 3 Problem Formulation: Optimizing Compound AI Systems
Forward execution. For a given input $x$ and configuration policy $\mathbf{v}$, the system executes components in topological order over the edge set $\mathcal{E}(x)$: $y_{k}=C_{k}\bigl(\{y_{i}\mid C_{i}\in pa(C_{k})\};\,\mathbf{v}_{k}\bigr),$ where $pa(C_{k})$ denotes all the parents of component $C_{k}$ over $\mathcal{E}(x)$. For clarity, we define the overall system as $f(x;\,\mathbf{v}):=y$, where $y$ is a collection of outputs from one or more components.
Optimization objective. Given a dataset $\mathcal{D}$ with initial inputs and a user-defined global reward function $R:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$ that evaluates the final system output, the optimization goal is to find the configuration policy $\mathbf{v}^{\star}(x)$ that maximizes the expected global reward:
@@ 4 Optimas: Globally Aligned Local Rewards for Optimization
Challenges. Directly optimizing the objective in Eq. 1 is difficult. As the configuration spaces are typically non-differentiable, gradient-based optimization cannot be used. Moreover, each policy $\mathbf{v}_{k}$ may control a different configuration type, so the joint policy $\mathbf{v}$ can span heterogeneous spaces. Therefore, previous efforts (Yuksekgonul et al., 2025; Zhao et al., 2025; Chen et al., 2025b; Khattab et al., 2023) largely focus on optimizing the policy for single types of configurations, which simplifies the optimization problem; however, this also leads to suboptimal compound systems.
Key intuition. To address the challenges, our approach (Figure 3) learns Local Reward Functions (LRFs) that align with the global reward for individual components, allowing local and independent optimization on heterogeneous components using different optimization approaches.
Such local-global alignments (Section 4.1) encourage that the global reward to increase during local optimizations (Section 4.3). Moreover, as the system configurations change during optimization, the LRFs should be adapted to remain aligned. To ensure alignment, Optimas employs a lightweight adaptation mechanism that updates LRFs with minimal data sampled from the system, preserving consistency with the global reward (Section 4.2).
@@ 4.1 Learning Local Reward Functions
Definition (Local Reward Function (LRF)). An LRF on component $C_{k}$ is defined as $r_{k}:(x_{k},y_{k})\rightarrow\mathbb{R}$, which evaluates the component’s output $y_{k}$ given the provided context $x_{k}$.
Implementation. We implement all LRFs with a LLM backbone $\phi$ and separate linear heads $h_{k}$ for a component $C_{k}$. The backbone encodes the concatenated text inputs $x_{k},y_{k}]$ into an embedding, and the corresponding head projects this embedding to a scalar reward value. Using such a multitask neural network ensures scalability with large number of components and reduces memory costs. Specifically, each LRF is modeled as:
$r_{k}(x_{k},y_{k})=h_{k}\circ\phi([x_{k},y_{k}])),\quad\text{for all }k\text{ if }\mathbf{v}_{k}\text{ is non-empty. }$ (2)
@@
where $\mathbf{v}_{-k}$ denotes the configurations of all downstream components (those that directly or indirectly receive information originating from $C_{k}$). The expected global reward for each candidate output is estimated via Monte Carlo sampling. This involves executing the downstream components with the candidate output fixed, capturing their stochasticity, and averaging the resulting global rewards from the final system outputs.
Objective of reward functions. To make each LRF $r_{k}$ conform to the local-global alignment property, we collect $\mathcal{D}_{k}(\mathbf{v})$, a preference dataset under the current system configuration $\mathbf{v}$, and train each $r_{k}$ using a pairwise log‑sigmoid ranking loss:
@@
The collection of $\mathcal{D}_{k}(\mathbf{v})$ follows the following steps: 1) execute the compound system up to $C_{k}$ and record the partial trajectory $\bigl\langle x,\,(x_{1},y_{1}),\dots,(x_{k-1},y_{\,k-1})\bigr\rangle$; 2) sample two candidate outputs for $C_{k}$ (e.g., via higher-temperature decoding or alternate hyperparameters); and 3) estimate their expected task metrics according to the expectation terms on the right-hand side of Eq. equation 3. The output with the higher expected value is labeled as $y_{k}^{+}$, and the other as $y_{k}^{-}$ in $\mathcal{D}_{k}(\mathbf{v})$.
4.2 Adaptive Local Reward Functions
Problem: Misaligned LRFs in the evolving system. As the system configuration changes during optimization, LRFs trained under a previous configuration $\mathbf{v}^{t}$ may become inaccurate under the updated configuration $\mathbf{v}^{t+1}$. Specifically, (1) after updating $C_{k}$, the same outputs from its upstream component $C_{i}$ ($i<k$) may lead to different global reward, making $r_{j}$ misaligned. (2) Its downstream components $C_{j}$ ($j>k$) now receive inputs generated by the updated $C_{k}$, which may fall outside the distribution seen by their LRFs. These shifts accumulate over time, degrading the local–global alignment property (Eq. 3) that LRFs are designed to satisfy.
However, retraining all LRFs from scratch after every configuration update is expensive. To address this, we develop a lightweight adaptation strategy that incrementally refines the LRFs as the system changes, maintaining alignment without full retraining.
Stage 1: Initial reward modeling. Given the initial system configuration and a dataset with initial inputs, we first construct an offline preference dataset for each component and train its LRF to convergence. This offline phase establishes well-aligned LRFs that accurately reflect each component’s contribution to the global reward.
Stage 2: Online reward function adaptation. We sample a small batch of input data and construct a mini-batch of preference data $\mathcal{B}_{k}$ for $C_{k}$ using the steps described in Section 4.1. We then optimize the LRF on $C_{k}$ on the objective $\mathcal{L}_{k}(\mathcal{B}_{k})$ following the definition in Eq. 4. To enable stable optimization and improve data efficiency, we maintain a buffer of previous generated preference data into $\mathcal{B}_{k}$ . This adaptation helps maintain the local–global alignment property in Eq. 3.